OpenStudy (mwhit2101):

How is 3/4 equal to 6/8 How is 6/8 equal to 12/16? Hewp pwz

OpenStudy (shamil98):

3/4 is a simplified form of 6/8 basically. \[\frac{ 3 }{ 4 } \times \frac{ 2 }{ 2 } = \frac{ 6 }{ 8 }\]

OpenStudy (mwhit2101):

I UNDERSTAND THANK YOU